### v1

# Enabling Automatic Differentiation with Mollified Graph Neural Operators

Abstract
Physics-informed neural operators offer a powerful framework for learning solution operators of partial differential equations (PDEs) by combining data and physics losses. However, these physics losses rely on derivatives. Computing these derivatives remains challenging, with spectral and finite difference methods introducing approximation errors due to finite resolution. Here, we propose the mollified graph neural operator ($m$GNO), the first method to leverage automatic differentiation and compute exact gradients on arbitrary geometries. This enhancement enables efficient training on irregular grids and varying geometries while allowing seamless evaluation of physics losses at randomly sampled points for improved generalization. For a PDE example on regular grids, $m$GNO paired with autograd reduced the L2 relative data error by 20× compared to finite differences, although training was slower. It can also solve PDEs on unstructured point clouds seamlessly, using physics losses only, at resolutions vastly lower than those needed for finite differences to be accurate enough. On these unstructured point clouds, $m$GNO leads to errors that are consistently 2 orders of magnitude lower than machine learning baselines (Meta-PDE) for comparable runtimes, and also delivers speedups from 1 to 3 orders of magnitude compared to the numerical solver for similar accuracy. $m$GNOs can also be used to solve inverse design and shape optimization problems on complex geometries.
Machine Learning, ICML
1 Introduction
Figure 1: Mollified GNO kernel’s neighborhood and weights versus those of the vanilla GNO kernel with differing point densities.
PDEs are critical for modeling physical phenomena relevant for scientific applications. Unfortunately, numerical solvers become very expensive computationally when used to simulate large-scale systems. To avoid these limitations, neural operators, a machine learning paradigm, have been proposed to learn solution operators of PDEs (Azizzadenesheli et al., 2024). Neural operators learn mappings between function spaces rather than finite-dimensional vector spaces, and approximate solution operators of PDE families (Li et al., 2020b; Kovachki et al., 2023). One example is the Fourier Neural Operator (FNO) (Li et al., 2020a), which relies on Fourier integral transforms with kernels parameterized by neural networks. Another example, the Graph Neural Operator (GNO) (Li et al., 2020b), implements kernel integration with graph structures and is applicable to complex geometries and irregular grids. The GNO has been combined with FNOs in the Geometry-Informed Neural Operator (GINO) (Li et al., 2023) to handle arbitrary geometries when solving PDEs. Neural operators have been successfully used to solve PDE problems with significant speedups (Li et al., 2021a; Kurth et al., 2022). They have shown great promise, primarily due to their ability to receive input functions at arbitrary discretizations and query output functions at arbitrary points, and also due to their universal operator approximation property (Kovachki et al., 2021).
Figure 2: Qualitative assessment of the smoothness of derivatives. We fit GINO with a PINO loss to the differentiable function $u(x,y)=\sin(4\pi xy)$ on $[0,1]^{2}$, and compare the analytic ground truth derivative $\partial_{xx}u(x,y)=-16\pi^{2}y^{2}\sin(4\pi xy)$ (left) with the numerical derivatives obtained using automatic differentiation (middle) and finite differences (right).
BurgersNonlinear PoissonHyperelasticity
Figure 3: Examples of solutions for the problems considered.
However, purely data-driven approaches may underperform in situations with limited or low resolution data (Li et al., 2021b), and may be supplemented using knowledge of physics laws (Karniadakis et al., 2021) as additional loss terms, as done in Physics-Informed Neural Networks (PINNs) (Raissi et al., 2017a, b, 2019). In the context of neural operators, the Physics-Informed Neural Operator (PINO) (Li et al., 2021b) combines training data (when available) with a PDE loss at a higher resolution, and can be finetuned on a given PDE instance using only the equation loss to provide a nearly-zero PDE error at all resolutions.
A major challenge when using physics losses is to efficiently compute derivatives without sacrificing accuracy, since numerical errors on the derivatives are compounded in the physics losses and output solutions. Approximate derivatives can be computed using finite differences (FDs), but can require a very high resolution grid to be sufficiently accurate, thus becoming intractable for fast-varying dynamics. Numerical derivatives can also be computed using Fourier differentiation, but this requires smoothness, uniform grids, and performs best when applied to periodic problems. In contrast, automatic differentiation computes exact derivatives using repeated applications of the chain rule and scales better to large-scale problems and fast-varying dynamics. Unlike numerical differentiation methods which introduce errors, automatic differentiation gives exact gradients, ensuring the accuracy required for physics losses, making it the preferred approach for physics-informed machine learning.
Approach.
 We propose a fully differentiable modification to GNOs to allow for the use automatic differentiation when computing derivatives and physics losses, which was previously prohibited by the GNO’s non-differentiability. More precisely, we replace the non-differentiable indicator function in the GNO kernel integration by a differentiable weighted function. This is inspired by mollifiers in functional analysis (Evans, 2010), which are used to approximate, regularize, or smooth functions. As a result, our differentiable mollified GNO ($m$GNO) is the first method capable of computing exact derivatives at arbitrary query points. The resulting $m$GNO can, in particular, be used within GINO to learn efficiently and accurately the solution operator of families of large-scale PDEs with varying geometries without data using physics losses.
We test the proposed approach on Burgers’ equation with regular grids, and nonlinear Poisson and hyperelasticity equations with varying domain geometries. Figure 3 displays examples of solutions, highlighting the complexity of the geometries considered). We show that physics losses are often critical, highlighting the need for efficient and accurate methods to compute derivatives. Using autograd instead of FDs leads to 20× reductions of the L2 relative data loss for Burgers’ equation on regular grids, suggesting that the Autograd physics loss better captures the physics underlying the data. Autograd $m$GINO performs seamlessly for the PDEs on unstructured point clouds, while FDs are not sufficiently accurate at the same training resolution and would need at least 9× more points to compute reasonable derivatives. Autograd $m$GINO achieves a relative error 2-3 orders of magnitude lower than the machine learning baselines (Meta-PDE (Qin et al., 2022) with LEAP and MAML) considered for a comparable running time, and enjoys speedups of 20-25× and 3000-4000× compared to the numerical solver for similar accuracy on the Poisson and hyperelasticity equations. Furthermore, as a result of its differentiability, $m$GINO can be used seamlessly for solving inverse design and shape optimization problems on complex geometries, as demonstrated with an airfoil design problem.
2 Background
2.1 Neural Operators
Neural operators compose linear integral operators $\mathcal{K}$ with pointwise non-linear activation functions $\sigma$ to approximate non-linear operators. A neural operator is defined as
$\mathcal{Q}\circ\sigma(W_{L}+\mathcal{K}_{L}+b_{L})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1}+b_{1})\circ\mathcal{P}$ (1)
where $\mathcal{Q}$ and $\mathcal{P}$ are the pointwise neural networks that encode (lift) the lower dimension function onto a higher dimensional space and project it back to the original space, respectively. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters of a FNO model consist of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. (2024) maintain a comprehensive open-source library for learning neural operators in PyTorch, which serves as the foundation for our implementation.
Fourier Neural Operator (FNO).
 A FNO (Li et al., 2020a) is a neural operator using Fourier integral operator layers
$\bigl{(}\mathcal{K}(\phi)v_{t}\bigr{)}(x)=\mathcal{F}^{-1}\Bigl{(}R_{\phi}\cdot(\mathcal{F}v_{t})\Bigr{)}(x)$ (2)
where $R_{\phi}$ is the Fourier transform of a periodic function $\kappa$ parameterized by $\phi$. On a uniform mesh, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform. The FNO architecture is displayed in Figure 8.
Graph Neural Operator (GNO).
 GNO (Li et al., 2020b) implements kernel integration with graph structures and is applicable to complex geometries and irregular grids. The GNO kernel integration shares similarities with the message-passing implementation of graph neural networks (GNN) (Battaglia et al., 2016). However, GNO defines the graph connection in a ball in physical space, while GNN assumes a fixed set of neighbors. The GNN nearest-neighbor connectivity violates discretization convergence and degenerates into a pointwise operator at high resolutions, leading to a poor approximation of the operator. In contrast, GNO adapts the graph based on points within a physical space, allowing for universal approximation of operators.
Specifically, the GNO acts on an input function $v$ as follows,
$\mathcal{G}_{\mathrm{GNO}}(v)(x)$ $\coloneqq\int_{D}\mathds{1}_{\text{B}_{r}(x)}(y)\kappa(x,y)v(y)\,\mathrm{d}y,$ (3)
where $D\subset\mathbb{R}^{d}$ is the domain of $v$, $\kappa$ is a learnable kernel function, and $\mathds{1}_{\text{B}_{r}(x)}$ is the indicator function over the ball $\text{B}_{r}(x)$ of radius $r>0$ centered at $x\in D$. The radius $r$ is a hyperparameter, and the integral can be approximated with a Riemann sum, for instance.
Geometry-Informed Neural Operator (GINO).
 GINO (Li et al., 2023) proposes to combine a FNO with GNOs to handle arbitrary geometries. More precisely, the input is passed through three main neural operators,
$\mathcal{G}_{\mathrm{GINO}}\ =\ \mathcal{G}_{\mathrm{GNO}}^{decoder}\ \circ\ \mathcal{G}_{\mathrm{FNO}}\ \circ\ \mathcal{G}_{\mathrm{GNO}}^{encoder}.$ (4)
First, a GNO encodes the input given on an arbitrary geometry into a latent space with a regular geometry. The encoded input can be concatenated with a signed distance function evaluated on the same grid if available. Then, a FNO is used as a mapping on that latent space for global integration. Finally, a GNO decodes the output of the FNO by projecting that latent space representation to the output geometry. The GINO architecture is displayed in Figure 9.
2.2 Physics-Informed Machine Learning
Physics-Informed Neural Network (PINN).
 In the context of solving PDEs, a PINN (Raissi et al., 2017a, b, 2019) is a neural network representation of the solution of a PDE, whose parameters are learned by minimizing the distance to the reference PDE solution and deviations from known physics laws such as conservation laws, symmetries and structural properties, and from the governing differential equations. PINNs minimize a composite loss
$\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ (5)
where $\mathcal{L}_{\text{data}}$ measures the error between data and model predictions wile $\mathcal{L}_{\text{physics}}$ penalizes deviations away from physics laws. PINN overcomes the need to choose a discretization grid of most numerical solvers, but only learns the solution for a single PDE instance and do not generalize to other instances without further optimization. Many modified versions of PINNs have also been proposed and used to solve PDEs in numerous contexts (Jagtap et al., 2020; Cai et al., 2022; Yu et al., 2022). When data is not available, we can try to learn the solution by minimizing $\mathcal{L}_{\text{physics}}$ only.
While physics losses can prove very useful, the resulting optimization task can be challenging and prone to numerical issues. The training loss typically has worse conditioning as it involves differential operators that can be ill-conditioned. In particular, Krishnapriyan et al. (2021) showed that the loss landscape becomes increasingly complex and harder to optimize as the physics loss coefficient $\lambda$ increases. The model could also converge to a trivial or non-desired solution that satisfies the physics laws on the set of points where the physics loss is computed (Leiteritz & Pflüger, 2021). There could also be conflicts between the multiple loss terms. Even without such conflicts, the losses can vary significantly in magnitude, leading to unbalanced back-propagated gradients during training (Wang et al., 2021). Manually tuning the loss coefficients can be computationally expensive, especially as the number of loss terms increases. Note that strategies have been developed to adaptively update the loss coefficients and mitigate this issue (Chen et al., 2018; Heydari et al., 2019; Bischof & Kraus, 2021).
Physics-Informed Neural Operator (PINO).
 In PINO (Li et al., 2021b), a FNO is trained with training data (when available) and physics losses at a higher resolution, allowing for near-perfect approximations of PDE solution operators. To further improve accuracy at test time, the trained model can be finetuned on the given PDE instance using only the equation loss and provides a nearly-zero error for that instance at all resolutions. PINO has been successfully applied to many PDEs (Song et al., 2022; Meng et al., 2023; Rosofsky et al., 2023). In practice, the PDE loss vastly improves generalization, physical validity, and data efficiency in operator learning compared to purely data-driven methods.
Physics-informed approaches without physics losses.
 Various approaches enforce physics laws in surrogate models without using physics losses. This can be achieved, for instance, by using projection layers (Jiang et al., 2020; Duruisseaux et al., 2024; Harder et al., 2024), by finding optimal linear combinations of learned basis functions that solve a PDE-constrained optimization problem (Negiar et al., 2022; Chalapathi et al., 2024), by leveraging known characterizations and properties of the solution operator as for divergence-free flows (Richter-Powell et al., 2022; Mohan et al., 2023; Xing et al., 2024), or Hamiltonian systems with their symplectic structure (Burby et al., 2020; Jin et al., 2020; Chen & Tao, 2021; Duruisseaux et al., 2023).
2.3 Computing Derivatives
In order to use physics losses, a major technical challenge is to efficiently compute derivatives without sacrificing accuracy, since numerical errors made on the derivatives will be amplified in the physics losses and output solution.
Finite Differences (FD).
 A simple approach is to use numerical derivatives computed using finite differences (FD). This differentiation method is fast and memory-efficient: given a $n$-points grid, it requires $O(n)$ computations. However, numerical differentiation using FD faces the same challenges as the corresponding numerical solvers: it requires a fine-resolution grid to be accurate and therefore becomes intractable for multi-scale and fast-varying dynamics. On point clouds, the stencil coefficients in FD formulas vary from point to point and must be computed each time, as outlined in Appendix G, adding to the computational cost. The errors in the resulting derivatives will also vary across the domain depending on the density of nearby points.
Fourier Differentiation.
 Fourier differentiation is also fast and memory-efficient to approximate derivatives as it requires $O(n\log n)$ given a $n$-points grid. However, just like spectral solvers, it requires smoothness, uniform grids, and performs best when applied to periodic problems. Fourier differentiation can be performed on non-uniform grids but the computational cost grows to $O(n^{2})$, and if the target function is non-periodic or non-smooth, the Fourier differentiation is not accurate. To deal with this issue, the Fourier continuation method (Maust et al., 2022) can be applied to embed the problem domain into a larger periodic space, at the cost of higher computational and memory complexity.
Pointwise Differentiation with Autograd.
 Derivatives can be computed pointwise using automatic differentiation by applying the chain rule to the sequence of operations in the model. Autograd (Maclaurin et al., 2015) automates this process by constructing a computational graph during the forward pass and leveraging reverse-mode differentiation to compute gradients during the backward pass. Autograd is typically the preferred method for computing derivatives in PINNs for a variety of reasons (Baydin et al., 2017): (1) Unlike FD, which introduces discretization errors, autograd provides exact derivatives (up to the limits of machine numerical precision), ensuring the accuracy that is critical in physics-informed machine learning, where derivative errors are amplified in physics losses. (2) Autograd computes gradients in a single pass (while FD requires multiple function evaluations). It provides exact derivatives regardless of the mesh resolution, making it particularly advantageous for large-scale problems, where FD become computationally intractable. (3) Autograd can compute higher-order derivatives with minimal additional cost, while FD require additional function evaluations and can suffer from further error accumulation. (4) Autograd can compute derivatives at any point in the domain seamlessly, while FD can struggle to handle complex geometries. (5) Figure 2 empirically shows that autograd derivatives are smoother and more stable than FD with the proposed model. However, autograd also has limitations: all operations need to be differentiable, and storing all intermediate computations in a computational graph can significantly increase memory usage for deep models. As a result, it can be slower and more memory-intensive than FD for simpler low-dimensional problems for which FD are accurate enough at low resolutions. When the physics loss involves a deep composition of operations, issues of vanishing or exploding gradients can also be exacerbated during backward propagation with automatic differentiation.
3 Methodology
Mollified GNO.
 Recall from Equation 3 that GNO computes the integral of an indicator function $\mathds{1}_{\text{B}_{r}(x)}$ that is not differentiable. We propose a fully differentiable layer that replaces the indicator function $\mathds{1}_{\text{B}_{r}(x)}$ with a differentiable weight function $w$ supported within $\mathds{1}_{\text{B}_{r}(x)}$. This is inspired by the mollifier in functional analysis (Evans, 2010), which is used to smooth out the indicator function. The resulting mollified GNO ($m$GNO) acts on an input function $v$ via
$\mathcal{G}_{m\mathrm{GNO}}(v)(x)\coloneqq\int_{\widetilde{D}}w(x,y)\kappa(x,y)v(y)\,\mathrm{d}y,$ (6)
for $x\in D$. Here, we padded the input of the FNO such that its output function $v$ is supported on the extended domain
$\widetilde{D}\coloneqq D+\text{B}_{r}(0)=\{x+y\ |\ x\in D,\ y\in\text{B}_{r}(0)\},$
which allows to recover exact derivatives at the boundary. Then, we can compute the derivative
$\partial_{x}\ \mathcal{G}_{m\mathrm{GNO}}(v)(x)\!=\!\!\int_{\text{B}_{r}(x)}{\!\!\!\!\partial_{x}\!\left[w(x,y)\kappa(x,y)\right]\!v(y)\,\mathrm{d}y}.$ (7)
Automatic differentiation algorithms can then be used to compute the derivatives appearing in physics losses. We can also use a cached version of neighbor search (i.e. store neighbors with nonzero weight) to keep the method as efficient as the original GNO up to the negligible cost of evaluating the weight function $w$. We emphasize that these definitions work for arbitrarily complex domains $D$.
As for the $m$GNO, the radius $r$ is a hyperparameter and the integral can be approximated with a Riemann sum, for instance. An example of simplified pseudocode for the $m$GNO layer is provided in Appendix C.
Mollified GINO.
The mollified GNO can then be used within a fully differentiable mollified GINO ($m$GINO) to learn efficiently the solution operator of PDEs with varying geometries using physics losses where the derivatives are computed using automatic differentiation,
$\mathcal{G}_{m\mathrm{GINO}}\ =\ \mathcal{G}_{m\mathrm{GNO}}^{decoder}\ \circ\ \mathcal{G}_{\mathrm{FNO}}\ \circ\ \mathcal{G}_{m\mathrm{GNO}}^{encoder}.$ (8)
This allows $m$GINO to be used for solving inverse design and shape optimization problems on complex geometries.
Weight Functions.
 Letting $d=\|x-y\|^{2}/r^{2}$, examples of weight functions $w$ with support in $\mathds{1}_{\text{B}_{r}(x)}(y)$ are given by
$w_{\text{bump}}(x,y)\coloneqq$ $\mathds{1}_{\text{B}_{r}(x)}(y)\ \exp\!\left(d^{2}/(d^{2}-1)\right),$ (9)
$w_{\text{quartic}}(x,y)\coloneqq$ $\mathds{1}_{\text{B}_{r}(x)}(y)\left(1-2d^{2}+d^{4}\right),$ (10)
$w_{\text{octic}}(x,y)\coloneqq$ $\mathds{1}_{\text{B}_{r}(x)}(y)\left(1-6d^{4}+8d^{6}-3d^{8}\right),$ (11)
$w_{\text{half\_cos}}(x,y)\coloneqq$ $\mathds{1}_{\text{B}_{r}(x)}(y)\left[0.5+0.5\cos{\!\left(\pi d\right)}\right].$ (12)
These weighting functions, displayed in Figure 10, are decreasing functions from $1$ to 00 on $[0,r]$.
4 Experiments
We use Meta-PDE (Qin et al., 2022) and the popular finite element method (FEM) FEniCS (Alnæs et al., 2015; Logg et al., 2011) as baselines. Meta-PDE learns initializations for PINNs over multiple instances that can be finetuned on any single instance, and two versions have been proposed based on the meta-learning algorithms MAML (Finn et al., 2017) and LEAP (Flennerhag et al., 2019). While FEniCS has been successfully applied in various disciplines, it has not been highly optimized for our applications and could possibly be outperformed by other FEMs and non-FEMs (Liu, 2009) such as Radial Basis Function (RBF) method (interpolation using RBFs), Finite Point Methods (FPMs) and Moving Least Squares (MLS) methods (weighted least squares to approximate solutions) and spectral methods (expanding the solution in a basis of functions).
4.1 Burgers’ Equation
Table 1: Relative L2 data loss and physics loss (computed using Autograd) for the proposed $m\mathrm{GNO}\circ\mathrm{FNO}$ models and PINO baselines, trained with $\mathcal{L}_{\text{Burgers}}$ where derivatives are computed using different methods.
Physics Loss Data Loss
Autograd (ours) $1.62\cdot 10^{-6}$ $\mathbf{1.33}\cdot\mathbf{10^{-2}}$
Finite Differences $4.89\cdot 10^{-3}$ $2.24\cdot 10^{-1}$
Fourier Differentiation $4.77\cdot 10^{-3}$ $2.19\cdot 10^{-1}$
PINO Baselines:
     Autograd $1.87\cdot 10^{-9}$ $1.22\cdot 10^{-1}$
     Finite Differences $3.71\cdot 10^{-4}$ $2.26\cdot 10^{-1}$
      Fourier Differentiation $3.63\cdot 10^{-4}$ $2.19\cdot 10^{-1}$
We consider the 1D time-dependent Burgers’ equation with periodic boundary conditions, whose description is provided in greater detail in Section D.1. We wish to learn the mapping $\mathcal{G}^{\dagger}$ from the initial condition $u(x,0)=u_{0}$ to the solution $u(x,t)$. Examples of solutions are shown in Figure 3. The input initial condition $u_{0}(x)$ given on a regular spatial grid is first duplicated along the temporal dimension to obtain a 2D regular grid, and then passed through a 2D FNO and a mollified GNO to produce a predicted function $v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ approximating the solution $u(x,t)$. We minimize a weighted sum $\mathcal{L}_{\mathrm{Burgers}}$ of the PDE residual and initial condition loss (see Equation 17).
We trained models using $\mathcal{L}_{\text{Burgers}}$, where the derivatives are computed using various differentiation methods. The results, obtained by finetuning the models one instance at a time and then averaging over the dataset, are presented in Table 1. Compared to numerical differentiation, $m\mathrm{GNO}\circ\mathrm{FNO}$ with Autograd was easier to tune and produced a model with data loss 20× lower (although 6× slower per training epoch), suggesting it better captures the physics underlying the data. It also achieved a physics loss 3000× smaller, but this was expected since these were evaluated using autograd.
A comparison to PINO (i.e. using a FNO instead of a $m\mathrm{GNO}\circ\mathrm{FNO})$ shows that PINO achieved a lower physics loss, but failed to reduce the Relative L2 data loss below $10\%$ on the hyperparameter sweep considered. Note that autograd $m\mathrm{GNO}\circ\mathrm{FNO}$ was only 1.2× slower than autograd PINO (although this was with a batch size of 1 and autograd PINO could be accelerated more easily using batching).
Table 2 displays the balance between training with physics and/or data losses. Using the data loss helps reduce the gap with reference data, but can come at then cost of a higher physics loss, and vice-versa.
We also tried pretraining the $m\mathrm{GNO}\circ\mathrm{FNO}$ models using the data loss before finetuning with $\mathcal{L}_{\text{Burgers}}$, but were not able to obtain better results this way. The data loss rapidly deteriorates from its original value obtained in a data-driven way, while the physics loss improves slowly but does not get better than the hybrid and physics only versions.
We also considered randomly subsampling the points at which the PDE residuals are evaluated, and Table 3 shows that despite the random location and reduction of the number of points used to evaluate the derivatives, we maintain good results, until the number of points becomes very low.
Table 2: Data loss (relative L2) and PDE loss of three models trained with different losses $\mathcal{L}$, using autograd to compute the physics losses, for the Burgers’ equation.
$\mathcal{L}_{\text{data}}$ $\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ $\mathcal{L}_{\text{physics}}$
Data $1.42\cdot 10^{-3}$ $3.81\cdot 10^{-3}$ $1.33\cdot 10^{-2}$
PDE $4.80\cdot 10^{-4}$ $1.48\cdot 10^{-4}$ $1.62\cdot 10^{-6}$
Table 3: Physics and data losses (relative L2) for models trained with the physics loss $\mathcal{L}_{\text{Burgers}}$ using autograd derivatives evaluated at different numbers of randomly subsampled points from the original $128\times 26$ regular grid of $3328$ points)
Physics Loss Data Loss
Full Grid (3328 points) $1.62\cdot 10^{-6}$ $1.33\cdot 10^{-2}$
Random 1000 points $6.70\cdot 10^{-6}$ $1.32\cdot 10^{-2}$
Random 500 points $8.15\cdot 10^{-6}$ $1.33\cdot 10^{-2}$
Random 250 points $1.47\cdot 10^{-5}$ $1.44\cdot 10^{-2}$
Random 100 points $4.06\cdot 10^{-5}$ $1.91\cdot 10^{-2}$
Random 50 points $5.31\cdot 10^{-5}$ $2.19\cdot 10^{-2}$
Random 25 points $1.29\cdot 10^{-4}$ $2.93\cdot 10^{-2}$
Random 10 points $3.49\cdot 10^{-4}$ $4.62\cdot 10^{-2}$
4.2 Nonlinear Poisson Equation
Figure 4: Computational time for inference versus accuracy for the nonlinear Poisson and hyperelasticity equations. We compare the proposed approach with Meta-PDE (MAML and LEAP) and with the baseline FEniCS solver. The FEniCS solver was used with 6 different resolutions for the Poisson equation, and 4 different resolutions (iteratively tries to refine the solution) for the hyperelasticity equation. The mean values across 200 instances are plotted, with the shaded regions representing the min and max error values.
We consider a nonlinear Poisson equation with varying source terms, boundary conditions, and geometric domain, whose description is detailed in Section D.2. Examples of solutions are shown in Figure 3. The mesh coordinates, signed distance functions, source terms, and boundary conditions, are passed through a GINO model of the form
$\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{\mathrm{GNO}}^{encoder}$
to approximate the solution $u$. We minimize a weighted sum $\mathcal{L}_{\mathrm{Poisson}}$ (21) of the PDE residual and boundary condition loss. In addition to the results of this section, an ablation study on the choice of GNO radius $r$ and weight function $w$ is provided in Appendix I.
4.2.1 Comparison to Baselines
The results in Figure 4 show the trade-off between inference time and accuracy. $m$GINO achieves a relative squared error 2-3 orders of magnitude lower than Meta-PDE for a comparable running time, and a speedup of 20-25× compared to the solver for similar relative accuracy. In addition, $m$GINO is more consistent across different instances, with smaller variations in errors compared to Meta-PDE.
Table 4: Data loss (MSE) and PDE loss of $m$GINO models trained with different losses $\mathcal{L}$ for Poisson and hyperelasticity equations.
$\mathcal{L}_{\text{data}}$ $\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ $\mathcal{L}_{\text{physics}}$
Poisson
Data $1.36\cdot 10^{-5}$ $2.29\cdot 10^{-5}$ $1.39\cdot 10^{-5}$
PDE $3.55\cdot 10^{2}$ $2.07\cdot 10^{-2}$ $7.21\cdot 10^{-3}$
 Hyperelasticity
Data $3.35\cdot 10^{-7}$ $9.69\cdot 10^{-7}$ $9.42\cdot 10^{-5}$
PDE $2.97\cdot 10^{26}$ $1.57\cdot 10^{-2}$ $1.21\cdot 10^{-2}$
(a)(b)(c)
Figure 5: (a) Prediction of a $m$GINO model trained using data loss only for the nonlinear Poisson equation. This is the predicted solution evaluated at a high resolution on a small patch centered at a latent query point of the $m$GNO. We see that the prediction exhibits discontinuities that coincide with the circles of radius $r$ (blue lines) centered at the neighboring latent query points. (b)(c) Norm of the gradient of the predicted solution shown in (a), computed using automatic differentiation (in (b)) and finite differences (in (c)).
4.2.2 Using Different Training Losses
We trained models using a data loss, a physics loss, and a hybrid loss (data and physics). Table 4 shows that using a physics loss has a comparable data error to the data-driven approach, while achieving a PDE residual 4-5 orders of magnitude lower. The hybrid approach proved more challenging to tune and did not perform as well.
When training only with data loss, the predicted solutions have discontinuities at higher resolutions coinciding with circles of radius $r$ centered at the latent query points of the GINO, (see Figure 5(a)). These discontinuities lead to high derivatives and inaccurate physics losses, regardless of the differentiation method used (see Figures 5(b)(c)). This happens despite the low data MSE, indicating that only training the $m$GINO model with data loss is not sufficient to capture the solution correctly at higher resolutions.
Recall that the GNO’s kernel integration can be viewed as an aggregation of messages if we construct a graph on the spatial domain of the PDE, as described in Li et al. (2020b).
The mean aggregation is given by
$v_{t+1}=\sigma\!\left(Wv_{t}(x)+\frac{1}{|N(x)|}\sum_{y\in N(x)}\kappa_{\phi}(e(x,y))v_{t}(y)\right)$
where $v_{t}(x)\in\mathbb{R}^{n}$ are the node features, $e(x,y)\in\mathbb{R}^{n_{e}}$ are the edge features, $W\in\mathbb{R}^{n\times n}$ is learnable, $N(x)$ is the neighborhood of $x$, and $\kappa_{\phi}(e(x,y))$ is a neural network mapping edge features to a matrix in $R^{n\times n}$. When using a differentiable weighting function, points in $N(x)$ at the edge of the neighborhood have near-zero weights but still contribute to the denominator $|N(x)|$. Thus, as the query point $x$ moves slightly, additional neighbors get included or excluded with near-zero weights, thereby introducing the discontinuities we see in Figure 5(a). Using a sum aggregation for the output GNO’s kernel integration mitigates the rigid patterns. When training with a physics loss, the patterns disappear, as shown in Figure 16, while the MSE for the predictions remains of the same order of magnitude.
Remark 4.1.
Finetuning the trained data-driven model using $\mathcal{L}_{\text{Poisson}}$ did not work well. The data-driven model has a physics loss 7 orders of magnitude larger than the data loss, and attempts at lowering the physics loss only worked by completely sacrificing the data loss.
4.2.3 Poisson Loss with Finite Differences (FD)
(a)(b)(c)
(d)(e)(f)
Figure 6: Prediction of a $m$GINO model and the norm of its gradient, computed on a $16\times 16$ regular grid (top row) and a $96\times 96$ regular grid (bottom row) using autograd and finite differences (FD). All plots are displayed on a $16\times 16$ grid for ease of qualitative comparison.
Figure 6 shows gradients norms on a domain patch, computed using autograd and FD at $16\times 16$ and higher resolutions (both displayed at $16\times 16$). The FD gradient norms at lower resolution in (c) differ completely from those obtained using autograd (b)(e) and high-resolution FD (f), while the latter are very similar. In addition, inaccuracies can be further amplified with FD on unstructured point clouds in regions with a low density of points. The autograd $m$GINOs were trained successfully with a density of randomly located points on the whole domain slightly lower than the density here with $16^{2}$ points on the domain patch displayed in Figure 6. Given that numerical errors made on derivatives are amplified in physics losses, $\mathcal{L}_{\text{Poisson}}$ does not provide enough information to move towards physically plausible solutions when computed using FD on unstructured point clouds at the same resolution at which autograd $m$GINOs models were trained successfully. This showcases how autograd $m$GINO can be used to obtain surrogate models using physics losses at resolutions at which FD is not accurate enough. We estimate that at least 9× more points would be necessary to compute reasonable FD derivatives. Higher resolution FD computations typically require a higher computational time and memory requirement (see Appendix H for a comparison of training times at different resolutions).
4.3 Hyperelasticity Equation
We consider a hyperelasticity equation modeling the deformation of a 2D porous hyperelastic material under compression with varying pore sizes, whose description is detailed in Section D.3. The mesh coordinates and signed distance functions are passed through a GINO model of the form
$\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{\mathrm{GNO}}^{encoder}$
to approximate the solution. This solution can be obtained as the minimizer of the total Helmholtz free energy of the system, which we use as our loss function instead of the PDE loss from the strong form of the hyperelasticity equation. We also add a weighted boundary loss term.
We trained models using data and/or physics losses. Table 4 shows that using a physics loss is critical for this problem. The data-driven approach achieves a low data loss, but the predictions do not satisfy the physics at all. In contrast, both the physics only and hybrid approaches achieve low data and PDE errors, with the hybrid approach performing the best. As for the Poisson equation, finetuning the trained data-driven model using the physics loss did not work, due to the high physics loss of the data-driven model. As in Section 4.2.3, autograd allowed to compute accurate derivatives at resolutions where finite differences are not accurate.
A comparison to baselines is displayed in Figure 4. $m$GINO achieves a relative squared error 2 orders of magnitude lower than Meta-PDE with a slightly faster running time, and achieves a speedup of 3000-4000× compared to the FEniCS solver for similar accuracy. In addition, $m$GINO achieves consistent results across different samples (all within a single order of magnitude) while Meta-PDE results span more than 3 orders of magnitude. This high variation in Meta-PDE results is likely caused by the difficulty disparity across samples, as samples with larger pores are harder to resolve.
Figure 7: The airfoil design problem. Parametrized vertical displacements of control spline nodes generate a point cloud, which is passed through the differentiable $m$GINO to obtain a pressure field $p$, from which we can compute the drag-lift ratio. We update the control nodes to minimize the drag-lift ratio by differentiating through this entire procedure.
4.4 Airfoil Inverse Design
We consider the transonic flow over an airfoil, governed by the Euler equation. An initial NACA-0012 shape is mapped onto a ‘cubic’ design element with $8$ control nodes in the vertical direction. That initial shape is morphed to a different shape following the displacement field of the control nodes. A more detailed description is provided in Section D.3
For the forward problem, the mesh point locations and signed distance functions are passed through a differentiable mollified GINO model of the form
$\mathcal{G}_{m\mathrm{GINO}}=\ \mathcal{G}_{m\mathrm{GNO}}^{decoder}\ \circ\ \mathcal{G}_{\mathrm{FNO}}\ \circ\ \mathcal{G}_{m\mathrm{GNO}}^{encoder}$ (13)
to produce an approximation of the pressure field $p$, achieving a validation (relative L2) data loss of $0.0146$ (trained in a data-driven manner). We use this trained model for inverse design. More precisely, we parametrize the shape of the airfoil by the vertical displacements of a few spline nodes, and set the design goal to minimize the drag-lift ratio. The parametrized displacements of the spline nodes are mapped to a mesh, which is passed through the $m$GINO to obtain a pressure field, from which we can obtain the drag lift ratio. We optimize the vertical displacement of spline nodes by differentiating through this entire procedure. A depiction of the airfoil design problem is given in Figure 7.
As a result of this optimization process, we obtain an airfoil design with drag coefficient 0.0216 and lift coefficient 0.2371, based on the model prediction. Using the numerical solver on this optimal design, we verify these predictions and obtain a similar drag coefficient 0.0217 and a similar lift coefficient 0.2532. This yields a drag-lift ratio of around 0.09, outperforming the optimal drag-lift ratio of 0.14 reported by Li et al. (2022) (drag 0.04 and lift 0.29, obtained using a Geo-FNO instead of $m$GINO).
Discussion
We proposed $m$GNO, a fully differentiable version of GNO, to allow for the use of automatic differentiation when computing derivatives, and embedded it within GINO to learn efficiently solution operators of families of large-scale PDEs with varying geometries without data. The proposed approach circumvents the computational limitations of traditional solvers, the heavy data requirement of fully data-driven approaches, and the generalization issues of PINNs.
The use of a physics loss proved critical in most experiments, and was sufficient to achieve good results in the absence of data. This highlights the need for efficient and accurate methods to compute physics losses, to improve data-efficiency and regularize neural operators. Autograd can compute exact derivatives in a single pass seamlessly across complex geometries and enables higher-order derivatives with minimal additional cost. However, it can be memory-intensive for deep models, possibly making it less efficient than finite differences (FDs) for simpler problems. Despite these limitations, its accuracy, capability to handle complex geometries, and scalability for complicated learning tasks make it the preferred differentiation method in PINNs, and a promising approach in our physics-informed neural operator setting on complex domains.
Using autograd instead of FDs led to a 20× reduction of the relative L2 data loss for Burgers’ equation on regular grids, suggesting that the autograd physics loss better captured the physics. underlying the data Autograd $m$GINO performed seamlessly for the Poisson and hyperelasticity equations on unstructured point clouds, while FDs were not sufficiently accurate at the training resolution used and would need at least 9× more points to compute reasonable derivatives. Autograd $m$GINO achieved a relative error 2-3 orders of magnitude lower than the Meta-PDE baselines for a comparable running time, and enjoyed speedups of 20-25× and 3000-4000× compared to the solver for similar accuracy on the Poisson and hyperelasticity equations. We demonstrated with an airfoil design problem that $m$GINO can be used seamlessly for solving inverse design and shape optimization problems on complex geometries.
In the future, we intend to investigate the use of adaptive loss balancing schemes for hybrid losses. Other mechanisms can improve the performance of PINNs, and we plan to explore their integration with the proposed approach. The training computational cost could also be amortized, for instance via higher-order automatic differentiation method such as the Stochastic Taylor Derivative Estimator (Shi et al., 2024).
Code
The PyTorch codes used for experiments presented in this paper have been added to the open-source neural operator library from Kossaifi et al. (2024) at https://github.com/neuraloperator/neuraloperator.
Acknowledgments
Ryan Lin is supported by the Caltech Summer Undergrad- uate Research Fellowships (SURF) program. David Pitt is supported by the Schmidt Scholars in Software Engineering program. Anima Anandkumar is supported in part by Bren endowed chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 senior fellow program at Schmidt Sciences.
Appendix A FNO and GINO architectures
Figure 8: The Fourier Neural Operator (FNO) architecture (extracted from (Li et al., 2020a)).
Figure 9: The Geometry-Informed Neural Operator (GINO) architecture (extracted from (Li et al., 2023)).
Appendix B Weighting Functions
Figure 10: Examples of weighting functions (9)-(12) for $m$GNO.
Appendix C $m$GNO Layer Pseudocode
Here, we provide a simplified example of PyTorch pseudocode for the $m$GNO layer
$\mathcal{G}_{m\mathrm{GNO}}(v)(x)\coloneqq\int_{\widetilde{D}}w(x,y)\kappa(x,y)v(y)\,\mathrm{d}y,$ (14)
with the half_cos weighting function
$w_{\text{half\_cos}}(x,y)=\mathds{1}_{\text{B}_{r}(x)}(y)\left[0.5+0.5\cos{\!\left(\pi d\right)}\right],$ (15)
where $d=\|x-y\|^{2}/r^{2}$.
Appendix D Problems Considered
D.1 Burgers’ Equation
The Burgers’ equation models the propagation of shock waves and the effects of viscosity in fluid dynamics. We consider the 1D time-dependent Burgers’ equation with periodic boundary conditions, and initial condition $u_{0}\in L_{\text{per}}^{2}(D;\mathbb{R})$ with $D=(0,1)$. The goal is to learn the mapping $\mathcal{G}^{\dagger}$ from the initial condition $u(x,0)=u_{0}$ to the solution $u(x,t)$ of the following differential equation for $x\in D$:
$\partial_{t}u(x,t)+\partial_{x}(u^{2}(x,t)/2)$ $=\nu\partial_{xx}u(x,t),\qquad t\in(0,1]$
$u(x,0)$ $=u_{0}(x).$
We focus on the dataset of Li et al. (2021a, b) consisting of 800 instances of the Burgers’ equation with viscosity coefficient $\nu=0.01$ and $128\times 26$ resolution. Each sample in the dataset corresponds to a different initial condition $u_{0}$ drawn from the Gaussian process $\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Examples of solutions are shown in Figure 11.
Figure 11: Ground truth solutions for several initial conditions of the Burgers’ Equation
The input initial condition $u_{0}(x)$ given on a regular spatial grid is first duplicated along the temporal dimension to obtain a 2D regular grid, and then passed through a 2D FNO and a mollified GNO to produce a predicted function
$v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ (16)
approximating the solution $u(x,t)$. We minimize a weighted sum of the PDE residual and initial condition loss,
$\mathcal{L}_{\mathrm{Burgers}}(v)\ =\ \left\|\partial_{t}v+\partial_{x}(v^{2}/2)-\nu\partial_{xx}v\right\|^{2}_{L^{2}(D\times(0,1))}\ +\ \alpha\left\|v(\cdot,0)-u_{0}\right\|^{2}_{L^{2}(D)}.$ (17)
D.2 Nonlinear Poisson Equation
The Poisson equation is a fundamental PDE that appears in numerous applications in science due to its ability to model phenomena with spatially varying and nonlinear behaviors. We consider the nonlinear Poisson equation with varying source terms, boundary conditions, and geometric domain,
$\nabla\cdot\left[(1+0.1u(\textbf{x})^{2})\nabla u(\textbf{x})\right]$ $=f(\textbf{x})$ $\textbf{x}\in\Omega$ (18)
$u(\textbf{x})$ $=b(\textbf{x})$ $\textbf{x}\in\partial\Omega$ (19)
where $u\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{2}$.
The domain $\Omega$ is centered at the origin and defined in polar coordinates with varying radius about the origin
$r(\theta)=r_{0}[1+c_{1}\cos(4\theta)+c_{2}\cos(8\theta)],$
where the parameters $c_{1}$ and $c_{2}$ are drawn from a uniform distribution on $(-0.2,0.2)$. The source term $f$ is a sum of radial basis functions $f(\textbf{x})=\sum_{i=1}^{3}\beta_{i}\exp{||\textbf{x}-\mu_{i}||_{2}^{2}},$ where $\beta_{i}\in\mathbb{R}$ and $\mu_{i}\in\mathbb{R}^{2}$ are both drawn from standard normal distributions. The boundary condition $b$ is a periodic function, defined in polar coordinates as $b_{0}+\frac{1}{4}[b_{1}\cos(\theta)+b_{2}\sin(\theta)+b_{3}\cos(2\theta)+b_{4}\sin(2\theta)]$, where the parameters $b_{i}$ are drawn from a uniform distribution on $(-1,1)$. This is the setting used by Qin et al. (2022).
The mesh coordinates, signed distance functions, source terms, and boundary conditions, are passed through a GINO model of the form
$\mathcal{G}_{m\mathrm{GNO}}^{decoder}\ \circ\ \mathcal{G}_{\mathrm{FNO}}\ \circ\ \mathcal{G}_{\mathrm{GNO}}^{encoder}$ (20)
to produce an approximation to the solution $u$. We minimize the PDE residual and boundary condition loss,
$\mathcal{L}_{\mathrm{Poisson}}(v)\ =\ \left\|\nabla\cdot\left[(1+0.1v^{2})\nabla v\right]-f\right\|^{2}_{L^{2}(\Omega)}\ +\ \alpha\left\|v-b\right\|^{2}_{L^{2}(\partial\Omega)}.$ (21)
Examples of predictions made by the trained $m$GINO model for a variety of geometries are displayed in Figure 12, below the corresponding reference solutions.
Figure 12: Comparison of solutions to nonlinear Poisson equations (top) with the corresponding $m$GINO predictions (bottom).
D.3 Hyperelasticity Equation
The hyperelasticity equation models the shape deformation under external forces of hyperelastic materials (e.g. rubber) for which the stress-strain relation is highly nonlinear. We consider the deformation of a homogeneous and isotropic hyperelastic material when compressed uni-axially (assuming no body and traction forces), and learn the final deformation displacement $u$ mapping the initial reference position to the deformed location. More precisely, we consider the deformation of a two-dimensional porous hyperelastic material under compression, as in Overvelde & Bertoldi (2014) and Qin et al. (2022). We keep the pores circular and fix the distance between the pore centers, so that the size of the pore is the only varied parameter. The size of the pores determines the porosity of the structure and affects the macroscopic deformation behavior of the structure. Examples of final deformation displacement fields are shown in Figure 13.
Figure 13: Final deformation displacement for several instances of the hyperelasticity dataset.
The mesh coordinates and signed distance functions are passed through a GINO model of the form
$\mathcal{G}_{m\mathrm{GNO}}^{decoder}\ \circ\ \mathcal{G}_{\mathrm{FNO}}\ \circ\ \mathcal{G}_{\mathrm{GNO}}^{encoder}$ (22)
to produce an approximation to the solution $u$ to the hyperelasticity equations.
[…]
D.4 Airfoil Inverse Design
We consider the transonic flow over an airfoil (ignoring the viscous effect), governed by the Euler equation,
$\frac{\partial\rho^{f}}{\partial t}+\nabla\cdot(\rho^{f}\mathbf{v})$ $=0,$ (23)
$\frac{\partial E}{\partial t}+\nabla\cdot\left((E+p)\mathbf{v}\right)$ $=0,$ (24)
$\frac{\partial\rho^{f}\mathbf{v}}{\partial t}+\nabla\cdot(\rho^{f}\mathbf{v}\otimes\mathbf{v}+p\mathbb{I})$ $=0,$ (25)
where $\rho^{f}$ is the fluid density, $\mathbf{v}$ is the velocity vector, $p$ is the pressure, and $E$ is the total energy. The far-field boundary condition is $\rho_{\infty}=p_{\infty}=1,\ M_{\infty}=0.8,\ AoA=0,$ where $M_{\infty}$ is the Mach number and $AoA$ is the angle of attack, and no-penetration condition is imposed at the airfoil.
We use the same dataset as Li et al. (2022), where the shape parameterization of the airfoil follows the design element approach (Farin, 2014). The initial NACA-0012 shape is mapped onto a ‘cubic’ design element with $8$ control nodes in the vertical direction with prior $d\sim\mathbb{U}[-0.05,0.05]$. That initial shape is morphed to a different shape following the displacement field of the control nodes.
The dataset contains $1000$ training samples and $200$ test samples generated using a second-order implicit finite volume solver. The C-grid mesh with ($220\times 50$) quadrilateral elements is used and adapted near the airfoil but not around the shock.
For the forward pass, the mesh point locations and signed distance functions are passed through a trained differentiable mollified GINO model of the form
$\mathcal{G}_{m\mathrm{GINO}}=\ \mathcal{G}_{m\mathrm{GNO}}^{decoder}\ \circ\ \mathcal{G}_{\mathrm{FNO}}\ \circ\ \mathcal{G}_{m\mathrm{GNO}}^{encoder}$ (26)
to produce an approximation of the pressure field $p$.
For the inverse problem, we parametrize the shape of the airfoil by the vertical displacements of a few spline nodes, and set the design goal to minimize the drag-lift ratio. The parametrized displacements of the spline nodes are mapped to a mesh, which is passed through the $m$GINO to obtain a pressure field, from which we can obtain the drag lift ratio. We optimize the vertical displacement of spline nodes by differentiating through this entire procedure.
A depiction of the airfoil design problem is given in Figure 7 and repeated below
Figure 14: The airfoil design problem. Parametrized vertical displacements of control spline nodes generate a point cloud, which is passed through the differentiable $m$GINO to obtain a pressure field $p$, from which we can compute the drag-lift ratio. We update the control nodes to minimize the drag-lift ratio by differentiating through this entire procedure.
Appendix E Choices of Hyperparameters
For all experiments, the “optimal” hyperparameters used (including weighting functions and loss coefficients) were obtained by conducting a grid search on a subset of hyperparameters. For the radius cutoff used in the weighting functions of $m$GNOs, we use the same radius as for the GNO to avoid the additional cost of carrying a second neighbor search.
We use the following Mean Squared Error (MSE), Relative Squared Error, L2 error, and Relative L2 errors, as metrics in our experiments:
$\textsc{MSE}(y_{\text{pred}},y_{\text{true}})=\frac{1}{N}\sum_{i}^{N}{\left(y_{\text{pred}}^{(i)}-y_{\text{true}}^{(i)}\right)^{2}},\qquad\textsc{RelativeSquaredError}(y_{\text{pred}},y_{\text{true}})=\frac{\sum_{i}^{N}{\left(y_{\text{pred}}^{(i)}-y_{\text{true}}^{(i)}\right)^{2}}}{\sum_{i}^{N}{{y_{\text{true}}^{(i)\ 2}}}\ +\ \varepsilon}$
[…]
where $\varepsilon$ is a small positive number for numerical stability, and the constant $C$ is a scaling constant taking into account the measure an dimensions of the data, to ensure that the loss is averaged correctly across the spatial dimensions.
E.1 Burgers’ Equation
For Burger’s equation, the input initial condition $u_{0}(x)$ given on a regular spatial grid in the domain $[0,1]^{2}$ is first duplicated along the temporal dimension to obtain a 2D regular grid of resolution $128\times 26$, and then passed through a 2D FNO and a mollified GNO to produce a predicted function $v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ approximating the solution $u(x,t)$.
For this experiment, the 2D FNO has $4$ layers, each with $26$ hidden channels and $(24,24)$ Fourier modes, and we used a a Tucker factorization with rank $0.6$ of the weights. The $m$GNO uses the half_cos weighting function with a radius of $0.1$, and a $2$-layer MLP with $[64,64]$ nodes. The resulting model has $1,019,569$ trainable parameters, and was trained in PyTorch for $10,\!000$ epochs using the Adam optimizer with learning rate $0.002$ and weight decay $10^{-6}$, and the ReduceLROnPlateau scheduler with factor $0.9$ and patience $50$.
E.2 Nonlinear Poisson Equation
For the nonlinear Poisson equation, the mesh coordinates within $[-1.4,1.4]^{2}$, signed distance functions, source terms, and boundary conditions, are passed through $\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{\mathrm{GNO}}^{encoder}$ model to produce an approximation to the solution $u$.
For this experiment, the input GNO has a radius of $0.16$, and a $3$-layer MLP with $[256,512,256]$ nodes. The 2D FNO has $4$ layers, each with $64$ hidden channels and $(20,20)$ Fourier modes, and acts on a latent space of resolution $64\times 64$. The output $m$GNO uses the half_cos weighting function with a radius of $0.175$, and a $3$-layer MLP with $[512,1024,512]$ nodes. The resulting model has $8,691,972$ trainable parameters, and was trained in PyTorch for $300$ epochs using the Adam optimizer with learning rate $0.0001$ and weight decay $10^{-6}$, and the ReduceLROnPlateau scheduler with factor $0.9$ and patience $2$. We used $7000$ samples for training and $3000$ samples for testing.
E.3 Hyperelasticity Equation
For the hyperelasticity equation, the mesh coordinates within $[0,1]^{2}$ and signed distance functions are passed through a $\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{\mathrm{GNO}}^{encoder}$ model to produce an approximation to the solution $u$ to the hyperelasticity equations.
For this experiment, the input GNO has a radius of $0.05625$, and a $3$-layer MLP with $[128,256,128]$ nodes. The 2D FNO has $4$ layers, each with $64$ hidden channels and $(20,20)$ Fourier modes, and acts on a latent space of resolution $32\times 32$. The output $m$GNO uses the half_cos weighting function with a radius of $0.1125$, and a $3$-layer MLP with $[1024,2048,1024]$ nodes. The resulting model has $11,678,211$ trainable parameters, and was trained in PyTorch for $400$ epochs using the Adam optimizer with learning rate $0.0001$ and weight decay $10^{-6}$, and the ReduceLROnPlateau scheduler with factor $0.9$ and patience $10$. We used $1000$ samples for training and $1000$ samples for testing.
E.4 Airfoil Inverse Design
For the airfoil design forward problem, the $220\times 50$ mesh point locations within $[-40,40]^{2}$ and signed distance functions are passed through a differentiable mollified GINO model of the form $\mathcal{G}_{m\mathrm{GINO}}=\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{m\mathrm{GNO}}^{encoder}$ to produce an approximation of the pressure field $p$.
For this experiment, the input $m$GNO uses the half_cos weighting function with a radius of $2$, and a $3$-layer MLP with $[128,128,128]$ nodes. The 2D FNO has $3$ layers, each with $16$ hidden channels and $(36,36)$ Fourier modes, and acts on a latent space of resolution $64\times 64$. The output $m$GNO uses the half_cos weighting function with a radius of $6$, and a $3$-layer MLP with $[128,128,128]$ nodes. The resulting model has $1,162,546$ trainable parameters, and was trained in PyTorch for $750$ epochs using the Adam optimizer with learning rate $0.0001$ and weight decay $10^{-9}$, and the ReduceLROnPlateau scheduler with factor $9$ and patience $5$. We used $1000$ samples for training and $200$ samples for testing.
Appendix F Effect of Using Different Training Losses
(a)(b)(c)
Figure 15: (a) Prediction of a $m$GINO model trained using data loss only for the nonlinear Poisson equation. This is the predicted solution evaluated at a high resolution on a small patch centered at a latent query point of the $m$GNO. We see that the prediction exhibits discontinuities that coincide with the circles of radius $r$ (blue lines) centered at the neighboring latent query points. (b)(c) Norm of the gradient of the predicted solution shown in (a), computed using automatic differentiation (in (b)) and finite differences (in (c)).
Figure 16: Comparison between the prediction of a $m$GINO model (left) trained using a hybrid loss, $\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ and the corresponding ground truth solution (right) for the nonlinear Poisson equation. The prediction and ground truth solutions are evaluated at a high resolution on a small patch centered at a latent query point of the $m$GNO. Using the physics loss regularized the higher resolution prediction, and removed the visible discontinuity patterns shown in Figure 15(a).
Appendix G Finite Differences on Point Clouds
On regular grid, standard well-known stencils formulas are available to compute first-order and higher-order derivatives using finite differences. However, on arbitrary point clouds, the varying distances between points have to be taken into account and a different stencil is needed for each point at which the derivatives need to be computed. We detail below one strategy to obtain these stencil formulas in 2D.
We consider the case where we have an arbitrary point 2D cloud with $N$ points, $\{(x_{i},y_{i})\}_{i=1}^{N}$, and suppose that the function values $\{f(x_{i},y_{i})\}_{i=1}^{N}$ of a function $f$ are known at these points. The goal is to approximate partial derivatives of $f$ at any other point $(\tilde{x},\tilde{y})$ in the domain. We start with first-order derivatives, and write them as a finite difference with unknown stencil coefficients:
$\frac{\partial f}{\partial x}(\tilde{x},\tilde{y})\approx\sum_{i=1}^{N}c_{i}^{(x)}f(x_{i},y_{i}),\qquad\quad\frac{\partial f}{\partial y}(\tilde{x},\tilde{y})\approx\sum_{i=1}^{N}c_{i}^{(y)}f(x_{i},y_{i}),$ (27)
To find the stencil coefficients $c_{i}^{(x)}$ and $c_{i}^{(y)}$, we enforce that the approximation holds exactly for the functions $1$, $x$ and $y$, that is, we enforce that the approximation holds true for any polynomial of degree 1 in 2D. This results in the following systems of equations for the stencil coefficients $c_{i}^{(x)}$ and $c_{i}^{(y)}$,
$\sum_{i=1}^{N}c_{i}^{(x)}=0,\qquad\sum_{i=1}^{N}c_{i}^{(x)}(x_{i}-\tilde{x})=1,\qquad\sum_{i=1}^{N}c_{i}^{(x)}(y_{i}-\tilde{y})=0,$
and
$\sum_{i=1}^{N}c_{i}^{(y)}=0,\qquad\sum_{i=1}^{N}c_{i}^{(y)}(x_{i}-\tilde{x})=0,\qquad\sum_{i=1}^{N}c_{i}^{(y)}(y_{i}-\tilde{y})=1.$
This can be written as
$A\mathbf{c}^{(x)}=\mathbf{b}^{(x)},\qquad\qquad A\mathbf{c}^{(y)}=\mathbf{b}^{(y)},$
where
$A=\begin{bmatrix}1&1&\cdots&1\\ x_{1}-\tilde{x}&x_{2}-\tilde{x}&\cdots&x_{N}-\tilde{x}\\ y_{1}-\tilde{y}&y_{2}-\tilde{y}&\cdots&y_{N}-\tilde{y}\end{bmatrix},$
and
$\mathbf{c}^{(x)}=\begin{bmatrix}c_{1}^{(x)}\\ ...\\ c_{N}^{(x)}\end{bmatrix},\qquad\quad\mathbf{c}^{(y)}=\begin{bmatrix}c_{1}^{(y)}\\ ...\\ c_{N}^{(y)}\end{bmatrix},\qquad\quad\mathbf{b}^{(x)}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},\qquad\quad\mathbf{b}^{(y)}=\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}.$
These systems of equations can be solved using least squares:
$\mathbf{c}^{(x)}=(A^{\top}\!A)^{-1}A^{\top}\mathbf{b}^{(x)},\qquad\qquad\mathbf{c}^{(y)}=(A^{\top}\!A)^{-1}A^{\top}\mathbf{b}^{(y)}.$
Plugging these coefficients in Equation 27 gives the desired approximation to the first-order derivatives at $f$.
We emphasize that this procedure needs to be repeated for every point $(\tilde{x},\tilde{y})$ at which the derivatives need to be evaluated since the location of the point affects the entries of the matrix $A$. Note that it is not necessary and not recommended to use all $N$ points to approximate the derivatives, and one should instead identify a subset of nearest neighbors from which the finite differences can be computed.
To compute higher-order derivatives, one could follow a similar strategy (which would lead to a more complicated system of equations to solve), or first evaluate first derivatives on the point cloud and repeat the above procedure by replacing $f$ by its appropriate partial derivative.
Appendix H Training times of 2D $m$GNOs with Autograd and Finite Differences (FD)
We compare the training times per epoch when training 2D $m$GNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. Figures 17 and 18 display the training times per epoch versus the latent resolution and output resolution, respectively.
As expected, we see from Figure 17 that autograd is more expensive than FD on a uniform grid at fixed latent and output resolutions. We can also see, by looking at individual columns corresponding to fixed latent space resolutions (i.e. the same model architecture), how FD at higher output resolutions compare to autograd at lower output resolutions. In particular, non-uniform finite differences are often more expensive than autograd at a slightly lower resolution. Figure 18 quantifies how running time increases when the latent space resolution of the model increases.
Figure 17: Time per epoch when training 2D $m$GNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. We display the training times per epoch at different latent space resolutions (x-axis) and different output resolutions (the numbers next to the points denote the numbers of points along each dimension).
Figure 18: Time per epoch when training 2D $m$GNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. We display the training times per epoch at different output resolutions (x-axis) and different latent space resolutions (the numbers at the end of the curves denote the numbers of points along each dimension).
Appendix I Ablation study on GNO radius and weighting function
We investigate how performance changes as the GNO radius changes for GINO models, and also compare the performance of the different weighting functions (9)-(12) on the nonlinear Poisson equation. Note that the radius cutoff used in the weighting functions of $m$GNOs is the same as the GNO radius to avoid the additional cost of carrying a second neighbor search. The results are displayed in Table 5.
The GNO radius $r$ is an important hyperparameter to tune. It needs to be large enough so that the ball $\text{B}_{r}(x)$ contains sufficiently many other latent query points, and a value too large will lead to prohibitive computational and memory costs. We denote the number of latent query points in $\text{B}_{r}(x)$ by #$\#|\text{B}_{r}(x)|$. For the nonlinear Poisson equation, we are using a regular 2D latent space grid, so $\text{B}_{r}(x)$ will only contain a single latent query point ($x$ itself) until $r$ is large enough for $\text{B}_{r}(x)$ to contain one extra latent query point in each direction. $\text{B}_{r}(x)$ will progressively contain more latent query points by thresholds as the radius increases further together with computational time and memory cost. Table 5 shows the results obtained with different weighting functions for values of $r$ such that #$\#|\text{B}_{r}(x)|=3^{2},5^{2},7^{2},9^{2},15^{2}$. When the radius is too small, performance is poor. On the other hand, when $r$ becomes too large, the computational and memory costs significantly increase while performance deteriorates. The best performance was achieved with a radius for which #$\#|\text{B}_{r}(x)|=7^{2}$, and $w_{\text{half\_cos}}$ and $w_{\text{quartic}}$ led to the best performance.
Table 5: Validation MSE of $m$GINO models, trained on only PDE loss, with different GNO radius values and weighting functions.
bump half_cos quartic octic
#$\#|\text{B}_{r}(x)|=9$         ($r=0.0875$)    $3.64\cdot 10^{-1}$    $3.41\cdot 10^{-1}$    $2.20\cdot 10^{-2}$    $3.48\cdot 10^{-1}$
#$\#|\text{B}_{r}(x)|=25$     ($r=0.13125$)    $4.51\cdot 10^{-5}$    $3.59\cdot 10^{-5}$    $3.95\cdot 10^{-5}$    $4.24\cdot 10^{-5}$
#$\#|\text{B}_{r}(x)|=49$          ($r=0.175$) $1.40\cdot 10^{-4}$ $\mathbf{1.39\cdot 10^{-5}}$ $2.93\cdot 10^{-5}$ $9.49\cdot 10^{-5}$
#$\#|\text{B}_{r}(x)|=81$     ($r=0.21875$)    $5.35\cdot 10^{-4}$    $1.01\cdot 10^{-4}$    $7.79\cdot 10^{-5}$    $1.76\cdot 10^{-4}$
#$\#|\text{B}_{r}(x)|=225$         ($r=0.35$) $3.92\cdot 10^{-3}$ $9.36\cdot 10^{-4}$ $4.28\cdot 10^{-3}$ $6.59\cdot 10^{-4}$